# An iterative CMB lensing estimator minimizing instrumental noise bias

### v1

@@ II.1 Maximum a posteriori lensing estimator
Here, the column vector ${X^{\rm dat}}(\hat{n})=(T^{\rm dat},Q^{\rm dat},U^{\rm dat})^{T}(\hat{n})$ is the observed CMB data vector for temperature and $Q$ and $U$ Stokes parameters in real space; $\tilde{X}_{\bm{l}}=(T_{\bm{l}},E_{\bm{l}},B_{\bm{l}})^{T}$ are the unlensed CMB fields in multipole space (with $E_{\bm{l}}$ and $B_{\bm{l}}$ the $E$- and $B$-mode polarization multipoles, respectively); $\mathcal{B}$ is the beam and transfer function of the instrument; and $n$ the noise in the maps. The lensing deflection field $\bm{\alpha}$ is related to the lensing potential with $\bm{\alpha}=\nabla\phi$ (we neglect the curl potential), and $\mathcal{D}_{\bm{\alpha}}$ is the map-synthesis and deflection operator that transforms the unlensed CMB modes in multipole space into the lensed Stokes parameters in real space.
@@
At fixed lensing deflections, the CMB follows a Gaussian likelihood $\mathcal{L}$. Assuming a Gaussian prior, $-2\mathcal{\ln}\mathcal{P}(\bm{\alpha})=\sum_{\bm{L}}\left|\phi_{\bm{L}}\right|^{2}/C_{\bm{L}}^{\phi\phi}+\text{const.}$., on the lensing field, we can write, up to a constant, the log-posterior of the deflection field as
@@
$\hskip 65.04034pt-\frac{1}{2}\sum_{\bm{L}}\frac{\left|\phi_{\bm{L}}\right|^{2}}{C_{\bm{L}}^{\phi\phi}}\,.$ (3)
@@ II.2 Cross-only iterative estimator
where we also rescaled the normalisation by $n/(n-1)$ such that its expectation value over realisations of the CMB fields is equivalent to the co-add gradient of Eq. 9. This cross-only gradient is similar to the cross-only QE of Madhavacheril et al. (2020), adapted to the MAP formalism with the inclusion of the lensing and delensing operators in the filtering.
In practice, this cross-only gradient is the only modification we bring to the standard MAP estimator. Thus, the filtering operations, as well as the Hessian matrix estimate (following the L-BFGS scheme as in Carron and Lewis (2017)) used in the iterations, are unchanged with respect to the standard (co-add) MAP estimator. We reconstruct the cross-only lensing field $\phi^{\rm MAP,\times}$, by modifying the delensalot [^2] https://github.com/NextGenCMB/delensalot pipeline to rely on the cross-only gradient (10) during the iterations, instead of the full co-add one (4).
@@ II.5 Power spectrum from splits
Assuming we can decompose the CMB maps into $n$ splits, it is possible to estimate $\phi^{\rm MAP,\times}$ from any subset of these $n$ splits. For example, we can take a subset $a$ that contains $k$ maps to estimate $\phi^{\rm MAP,\times}_{a}$, and the complementary subset $b$ that contains the $n-k$ other maps to estimate $\phi^{\rm MAP,\times}_{b}$. The power spectrum $C_{L}(\phi^{\rm MAP,\times}_{a},\phi^{\rm MAP,\times}_{b})$ will avoid repetitions in the pairs of maps.
It follows that, if we have $n$ splits, we can estimate the lensing power spectrum with any combinations of the two subsets of maps of size $k$ and $n-k$, for all $k$.
$C_{L}^{\rm MAP,\times}=\frac{1}{N_{\rm sets}}\sum_{a,b}C_{L}(\phi^{\rm MAP,\times}_{a},\phi^{\rm MAP,\times}_{b})\,,$ (19)
@@
$C_{L}^{\rm MAP,\times}=\frac{1}{3}\left[C_{L}(\phi^{\rm MAP,\times}_{12},\phi^{\rm MAP,\times}_{34})\right.\\ +C_{L}(\phi^{\rm MAP,\times}_{13},\phi^{\rm MAP,\times}_{24})+\left.C_{L}(\phi^{\rm MAP,\times}_{14},\phi^{\rm MAP,\times}_{23})\right]\,,$ (20)
@@
For the cross-only MAP power spectrum, the disconnected contractions in the cross-power spectrum $C_{L}(\phi^{\rm MAP,\times}_{12},\phi^{\rm MAP,\times}_{34})$ that contributes to the $N_{L}^{(0)}$ bias, such as $X_{1}X_{3}$ and $X_{2}X_{4}$, are, by construction, free of noise contributions, since they avoid any auto-correlations between the split maps. In Section III.1 we will confirm with noiseless simulations that the cross-only MAP $N_{L}^{(0),\times}$ bias is indeed free from noise contributions.
@@ III.1 Isotropic noise
The noise bias is reduced by a factor of around five for the CMB-S4 polarization case. The reduction is less significant in the SO temperature case. This is expected as the CMB temperature is totally signal dominated in the range of multipoles we considered for the SO configuration (up to $\ell=3000$), so the bias reduction from cancellation of the noise auto-correlations is negligible. We see that there is a good agreement between the predicted and estimated noise biases for both the MAP and cross-only MAP estimators. The prediction for cross-only MAP with the CMB-S4 configuration is around $10\%$ lower than the simulation results. In a realistic analysis, this small offset would be captured by a realisation-dependent bias subtraction ($\mathrm{RD{\text{-}}}N_{L}^{(0)}$) Hanson et al. (2011), and by a Monte Carlo correction applied on the lensing bandpowers, but neither are applied here.
@@
Even if the biases are reduced with the cross-only MAP estimator, this does not correspond to a reduction in the variance of the lensing bandpowers over our set of simulations. We show in Fig. 2 the ratio of the bandpower errors for the cross estimators over the co-add estimators. In the temperature reconstruction, the increase in errors is negligible, in particular for the SO-like configuration, which is signal dominated. In the CMB-S4-like polarization case, the errors for the cross estimators are around 10% larger than the co-adds at high $L$, but below $L=1000$ (which in practice achieves most of the signal-to-noise ratio) the increase is below $5\%$ and so negligible.
Figure 2: Ratio of the CMB lensing power spectra variance, for the cross estimators over the standard (co-add) estimators, obtained from 200 full-sky simulations and binned in bandpowers. We show the QE (dashed lines) and MAP (solid lines) for our three experimental configurations. For the SO-like case we show both the temperature only and the temperature and polarization estimators. The variance of the cross estimators are always higher than the co-add estimators, as expected. This increase is, however, negligible for the SO temperature reconstruction since this is signal-dominated. We also see that in temperature, the increase in bandpower variance for the cross MAP estimator is similar to the cross QE estimator. For the CMB-S4 polarization configuration, the increases in variance for the cross estimators is a little worse, particularly for the MAP estimator. However, for the CMB-S4 case, at multipoles $L<1000$, where most of the signal-to-noise is accumulated, the degradation is still negligible.
@@ III.2 Anisotropic noise and mean-field
The left panel of Fig. 5 shows the mean-field for the ACT-like temperature configuration. The QE and MAP mean-fields are very similar since in this configuration the iterative estimator is almost equivalent to the QE. However, we see that the cross-only MAP reduces the mean-field power by about an order of magnitude. This mean-field is not totally nulled, contrary to the cross-only QE. This is likely due to the fact that during the iterative process, there are some non-trivial combinations of maps involved, as described schematically in Eq. 17.
[…]
Figure 5: Power spectra of the mean-field for the QE (orange), standard MAP (blue) and cross-only MAP (teal), estimated from averaging over 100 simulations with anisotropic noise patterns. The QE cross-only mean-field is zero by construction. For reference we show the fiducial lensing power spectrum (solid black line) and the $N_{L}^{(0),\rm MAP}$ bias (dashed black line). The left panel is for the ACT-like configuration with temperature, and the right panel is for CMB-S4 with polarization. We see that the cross-only MAP is able to reduce the mean-field term, in particular in the polarization S4 configuration. The mean field does not disappear in the ACT temperature configuration, but it is reduced by an order of magnitude.
@@ Appendix A Mean-field of the cross-only estimator
We derive here the expression for the additional mean-field term, $-\text{Tr}[\textrm{Cov}_{\bm{\alpha}}^{-1}N]/2$, that we introduce in the loss-function (11) for our split-only estimator. Assuming we split the data into $n$ split maps, we rewrite the block matrix $K_{\bm{\alpha}}^{-1}$ from Eq. (13) as a Kronecker product
@@
$-\frac{1}{2}{X^{\rm dat}_{\times}}^{T}K_{\times}^{-1}{X^{\rm dat}_{\times}}\right>_{\tilde{\bm{\alpha}}}&=-\frac{1}{2}\text{Tr}\left[K_{\times}^{-1}\widetilde{\textrm{Cov}}({X^{\rm dat}_{\times}})\right]\\ &=-\frac{1}{2}\text{Tr}\left[\left(\frac{1}{n(n-1)}(J_{n}-I_{n})\otimes\textrm{Cov}_{\bm{\alpha}}^{-1}\right)\left(J_{n}\otimes(\widetilde{\textrm{Cov}_{\bm{\alpha}}}-N)+I_{n}\otimes N_{\times}\right)\right]\\ &=-\frac{1}{2}\frac{1}{n(n-1)}\text{Tr}\left[(J_{n}-I_{n})J_{n}\otimes\textrm{Cov}_{\bm{\alpha}}^{-1}(\widetilde{\textrm{Cov}_{\bm{\alpha}}}-N)+(J_{n}-I_{n})\otimes\textrm{Cov}_{\bm{\alpha}}^{-1}N_{\times}\right]\\ &=-\frac{1}{2}\frac{1}{n(n-1)}\text{Tr}\left[(n-1)J_{n}\otimes(\textrm{Cov}_{\bm{\alpha}}^{-1}\widetilde{\textrm{Cov}_{\bm{\alpha}}}-\textrm{Cov}_{\bm{\alpha}}^{-1}N)\right]\\ &=-\frac{1}{2}\text{Tr}\left[\textrm{Cov}_{\bm{\alpha}}^{-1}\widetilde{\textrm{Cov}_{\bm{\alpha}}}\right]+\frac{1}{2}\text{Tr}\left[\textrm{Cov}_{\bm{\alpha}}^{-1}N\right]\,.\end{split}$ (27)
Taking the gradient of this expectation with respect to $\bm{\alpha}$, and then setting $\tilde{\bm{\alpha}}=\bm{\alpha}$, is equivalent to taking the gradient of $-{X^{\rm dat}_{\times}}^{T}K_{\times}^{-1}{X^{\rm dat}_{\times}}/2$ first and then averaging over data lensed by fixed $\bm{\alpha}$. Applied to the right of Eq. (27), this gives
$-\frac{1}{2}\text{Tr}\left[(\nabla_{\bm{\alpha}}\textrm{Cov}_{\bm{\alpha}}^{-1})\textrm{Cov}_{\bm{\alpha}}\right]+\frac{1}{2}\text{Tr}\left[(\nabla_{\bm{\alpha}}\textrm{Cov}_{\bm{\alpha}}^{-1})N\right]\\ =\frac{1}{2}\nabla_{\bm{\alpha}}(\ln\det\textrm{Cov}_{\bm{\alpha}})+\frac{1}{2}\nabla_{\bm{\alpha}}\text{Tr}\left[\textrm{Cov}_{\bm{\alpha}}^{-1}N\right]\,.$ (28)

### v2

@@ II.1 Maximum a posteriori lensing estimator
[…]
@@
At fixed lensing deflections, the CMB follows a Gaussian likelihood $\mathcal{L}$. Assuming a Gaussian prior, $2\mathcal{\ln}\mathcal{P}(\bm{\alpha})=\sum_{\bm{L}}\left|\phi_{\bm{L}}\right|^{2}/C_{\bm{L}}^{\phi\phi}+\text{const.}$, on the lensing field, we can write, up to a constant, the log-posterior of the deflection field as
@@
$\hskip 51.7479pt-\frac{1}{2}\sum_{\bm{L}}\frac{\left|\phi_{\bm{L}}\right|^{2}}{C_{\bm{L}}^{\phi\phi}}\,.$ (3)
@@ II.2 Cross-only iterative estimator
where we also rescaled the normalisation by n/(n1)n/(n-1) such that its expectation value over realisations of the CMB fields is equivalent to the co-add gradient of Eq. 9. This cross-only gradient is similar to the cross-only QE of Madhavacheril et al. (2020), adapted to the MAP formalism with the inclusion of the lensing and delensing operators in the filtering.
In practice, this cross-only gradient is the only modification we make to the standard MAP estimator. Thus, the filtering operations, as well as the Hessian matrix estimate (following the L-BFGS scheme as in Carron and Lewis (2017)) used in the iterations, are unchanged with respect to the standard (co-add) MAP estimator. We reconstruct the cross-only lensing field $\phi^{\rm MAP,\times}$, by modifying the delensalot [^2] https://github.com/NextGenCMB/delensalot pipeline to rely on the cross-only gradient (10) during the iterations, instead of the full co-add one (4).
@@ II.5 Power spectrum from splits
[…]
It follows that, if we have $n$ splits, we can estimate the lensing power spectrum with any combinations of the two subsets of maps of size $k$ and nkn-k, for all $k$.
[…]
@@
[…]
@@
[…]
@@ III.1 Isotropic noise
The noise bias is reduced by a factor of around five for the CMB-S4 polarization case. The reduction is less significant in the SO temperature case. This is expected as the CMB temperature is totally signal dominated in the range of multipoles we considered for the SO configuration (up to $\ell=3000$), so the bias reduction from cancellation of the noise auto-correlations is negligible. We see that there is a good agreement between the predicted and estimated noise biases for both the MAP and cross-only MAP estimators. The prediction for cross-only MAP with the CMB-S4 configuration is around $10\%$ lower than the simulation results for $L\sim 1500$, and around $50\%$ lower for $L\sim 200$. In a realistic analysis, this small offset would be captured by a realisation-dependent bias subtraction ($\mathrm{RD{\text{-}}}N_{L}^{(0)}$) Hanson et al. (2011), and by a Monte Carlo correction applied on the lensing bandpowers, but neither are applied here.
@@
Even if the biases are reduced with the cross-only MAP estimator, this does not correspond to a reduction in the variance of the lensing bandpowers over our set of simulations. We compare in Fig. 2 the bandpower errors for the cross-only estimators over the standard co-add MAP estimator. We see that the MAP cross-only estimator has a lower increase in variance than the cross-only QE estimator, in particular for the CMB-S4 polarization and the SO minimum-variance cases. But even for the ACT-like temperature configuration, the increase is reduced by a factor two compared to that for the cross-only QE. With the cross-only MAP in the temperature reconstruction, the increase in errors compared to the co-add MAP is negligible. For the minimum-variance SO-like configuration, the increase in variance of the cross-only MAP bandowers compared to the co-add MAP ones is below $5\%$ for all multipoles. In the CMB-S4-like polarization case, the errors for the cross-only MAP estimator are around 10% larger than the co-adds at high $L$, but below $L=1000$ (which in practice achieves most of the signal-to-noise ratio) the increase is below $5\%$ and so negligible.
Figure 2: Ratio of the CMB lensing power spectrum variance, for the cross-only estimators over the standard (co-add) MAP estimator, obtained from 200 full-sky simulations and binned in bandpowers. We show the cross-only QE (dashed lines) and cross-only MAP (solid lines), over the co-add MAP, for our three experimental configurations. For the SO-like case we show both the temperature-only and the temperature and polarization estimators. The variance of the cross-only estimators are always higher than the co-add MAP estimator, as expected. We see that the increase in variance of the cross-only QE is greatly reduced for the cross-only MAP estimator, in particular for surveys including polarization. Comparing the cross-only MAP to the co-add MAP, we see that the increase in variance is below $5\%$ for the ACT-like and SO-like configurations. For the CMB-S4 case, at multipoles $L<1000$, where most of the signal-to-noise is accumulated, the degradation is still negligible.
@@ III.2 Anisotropic noise and mean-field
In the left and right panels of Fig. 5 we show the mean-field for the ACT-like and SO-like temperature configurations. The QE and MAP mean-fields are very similar since in these configurations the iterative estimator is almost equivalent to the QE. However, we see that the cross-only MAP reduces the mean-field power by about an order of magnitude in both cases. This mean-field is not totally nulled, contrary to the cross-only QE. This is likely due to the fact that during the iterative process, there are some non-trivial combinations of maps involved, as described schematically in Eq. 17.
We show as dotted lines the predicted mean-field power spectra arising from noise anisotropies for both the standard QE and MAP estimators. These predictions are obtained using the mean-field (MF) response derived in Appendix B of Aghanim et al. (2020), generalized here to include the polarization-based estimator. For the MAP estimator, the MF response is computed using delensed CMB spectra (obtained via the iterative procedure employed in the calculation of the $N_{L}^{(0)}$ bias) in place of the lensed spectra. These predictions are found to be in good agreement with the mean-field power spectra estimated from simulations.
Figure 4: One realization of the noise map used for the simulations, following the Planck scanning strategy, scaled such that the its standard deviation is of $10\,\mu\text{K-arcmin}$.
Figure 5: Power spectra of the mean-field for the QE (orange), standard MAP (blue) and cross-only MAP (teal), estimated from averaging over 100 simulations with anisotropic noise patterns. The QE cross-only mean-field is zero by construction. For reference we show the fiducial lensing power spectrum (solid black line) and the $N_{L}^{(0),\rm MAP}$ bias (dashed black line). The left panel is for the ACT-like configuration with temperature, the central panel is for a SO-like configuration with temperature, and the right panel is for CMB-S4 with polarization. We see that the cross-only MAP is able to reduce the mean-field term, in particular in the polarization CMB-S4 configuration. The mean-field does not disappear in the temperature only reconstructions, but it is reduced by an order of magnitude for both ACT and SO. We also show the predicted mean-field power spectrum due to noise anisotropies as the dotted and dashed lines, for standard QE and MAP estimators, respectively.
@@ Appendix A Mean-field of the cross-only estimator
We derive here the expression for the additional mean-field term, $\text{Tr}[\textrm{Cov}_{\bm{\alpha}}^{-1}N]/2$, that we introduce in the loss-function (11) for our split-only estimator. Assuming we split the data into $n$ split maps, we rewrite the block matrix $K_{\bm{\alpha}}^{-1}$ from Eq. (13) as a Kronecker product
@@
$\begin{split}\left<-\frac{1}{2}{X^{\rm dat}_{\times}}^{T}K_{\times}^{-1}{X^{\rm dat}_{\times}}\right>_{\tilde{\bm{\alpha}}}&=-\frac{1}{2}\text{Tr}\left[K_{\times}^{-1}\widetilde{\textrm{Cov}}({X^{\rm dat}_{\times}})\right]\\ &=-\frac{1}{2}\text{Tr}\left[\left(\frac{1}{n(n-1)}(J_{n}-I_{n})\otimes\textrm{Cov}_{\bm{\alpha}}^{-1}\right)\left(J_{n}\otimes(\widetilde{\textrm{Cov}_{\bm{\alpha}}}-N)+I_{n}\otimes N_{\times}\right)\right]\\ &=-\frac{1}{2}\frac{1}{n(n-1)}\text{Tr}\left[(J_{n}-I_{n})J_{n}\otimes\textrm{Cov}_{\bm{\alpha}}^{-1}(\widetilde{\textrm{Cov}_{\bm{\alpha}}}-N)+(J_{n}-I_{n})\otimes\textrm{Cov}_{\bm{\alpha}}^{-1}N_{\times}\right]\\ &=-\frac{1}{2}\frac{1}{n(n-1)}\text{Tr}\left[(n-1)J_{n}\otimes(\textrm{Cov}_{\bm{\alpha}}^{-1}\widetilde{\textrm{Cov}_{\bm{\alpha}}}-\textrm{Cov}_{\bm{\alpha}}^{-1}N)\right]\\ &=-\frac{1}{2}\text{Tr}\left[\textrm{Cov}_{\bm{\alpha}}^{-1}\widetilde{\textrm{Cov}_{\bm{\alpha}}}\right]+\frac{1}{2}\text{Tr}\left[\textrm{Cov}_{\bm{\alpha}}^{-1}N\right]\,.\end{split}$ (27)
Taking the gradient of this expectation with respect to $\bm{\alpha}$, and then setting $\tilde{\bm{\alpha}}=\bm{\alpha}$, is equivalent to taking the gradient of ${X^{\rm dat}_{\times}}^{T}K_{\times}^{-1}{X^{\rm dat}_{\times}}/2$ first and then averaging over data lensed by fixed $\bm{\alpha}$. Applied to the right of Eq. (27), this gives
$\frac{1}{2}\text{Tr}\left[(\nabla_{\bm{\alpha}}\textrm{Cov}_{\bm{\alpha}}^{-1})\textrm{Cov}_{\bm{\alpha}}\right]+\frac{1}{2}\text{Tr}\left[(\nabla_{\bm{\alpha}}\textrm{Cov}_{\bm{\alpha}}^{-1})N\right]\\ =\frac{1}{2}\nabla_{\bm{\alpha}}(\ln\det\textrm{Cov}_{\bm{\alpha}})+\frac{1}{2}\nabla_{\bm{\alpha}}\text{Tr}\left[\textrm{Cov}_{\bm{\alpha}}^{-1}N\right]\,.$ (28)